askew616:

Simplify cube root of 7 over fifth root of 7. Will you guys help me.

askew616:

Hello

askew616:

Anybody know this

askew616:

Help

askew616:

Anybody?

askew616:

I know it is a hard question.

askew616:

Would you guys like answer choices?

askew616:

Can I get someone to reply?

mhanifa:

Can you right this down as exponents?

askew616:

Here are the choices 7 to the power of 1 over 5 7 to the power of 8 over 15 7 to the power of 5 over 3 7 to the power of 2 over 15

mhanifa:

This is not my question

mhanifa:

How would you show the cube root of 7?

askew616:

cube root of 7 over fifth root of 7

askew616:

That is the best I can show you. Thanks for trying to help.

mhanifa:

\[7^{1/3}\]

mhanifa:

What about the 5th root?

askew616:

Yor right here is the right question Simplify cube root of 7 over fifth root of 7

mhanifa:

\[7^{1/3}/7^{1/5}\]

mhanifa:

You need to use the following rule:

mhanifa:

\[a^b/a^c=a^{b-c}\]

mhanifa:

Can you apply this to your question?

askew616:

I don't understand

mhanifa:

\[7^{1/3}/7^{1/5}=7^{1/3-1/5}\] Can you find the exponent

askew616:

One question are most users on here kids?

mhanifa:

What is 1/3 - 1/5?

askew616:

What?

askew616:

I sorry I guess I am wasting your time.

mhanifa:

1/3 - 1/5 = 5/15 - 3/15= 2/15

askew616:

I meant I'm sorry

mhanifa:

Can you guess the choice?

askew616:

What do you think the answer is out of the choices I gave u

askew616:

I think it is D.

askew616:

What do you think?

mhanifa:

I gave you the solution and you guessed it right

askew616:

Thanks a lot! I will fan you.

mhanifa:

You are welcome!